OpenStudy (anonymous):

What is the solution to the rational equation 5 over x¨Csquared plus 3x plus 2 - 1 over x plus 2 = 1 over 3x plus 3 ? x = 7 x = 5 over 2 x = 5 x = 4

OpenStudy (anonymous):

can you draw the problem please ?

jhonyy9 (jhonyy9):

5 1 1 ------------ - --------- = --------- x^2 +3x +2 x +2 3x +3 this is correct ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so than x^2 +3x +2 = (x +1)(x+2) 3x+3 = 3(x+1) 5 1 1 ------------ - --------- = --------- x^2 +3x +2 x +2 3x +3 5 1 1 ---------- - ------- = -------- (x+1)(x+2) x+2 3(x+1) so than the common denominator will be 3(x+1)(x+2) 3*5 - 3(x+1) = x+2 15 - 3x -3 = x+2 12 -3x = x+2 12 -2 = x +3x 10 = 4x x = 10/4 x= 5/2

OpenStudy (anonymous):

thanks :)

jhonyy9 (jhonyy9):

was my pleasure yw good luck